OpenStudy (anonymous):

If a number n is subtracted from 18, the result is four less than n. What is the value of n?

OpenStudy (johnweldon1993):

So set up the equation 18 - n = n - 4 And solve for 'n'

OpenStudy (anonymous):

This will be =>18-n=n-4 =>2n=22 =>n=11

OpenStudy (anonymous):

Do I Subtract 18 From Both Sides?

OpenStudy (anonymous):

Always take all variables on one side & all constants on other side while solving such type of questions . (Here n is variable to be found out & the numbers are constants )

OpenStudy (johnweldon1993):

Right...so you would add 'n' to both sides You'll then have 18 = 2n - 4 Then you want to isolate 2n...so add 4 to both sides 22 = 2n And you want to solve for 'n' so divide both sides by 2 11 = n